OpenStudy (anonymous):

(x + 4)^2 + 4x = 0 Put the equation below in the form ax2 + bx + c = 0 :I

OpenStudy (accessdenied):

multiply out the binomial (x+4)^2 and combine like terms

OpenStudy (anonymous):

\[(x+4)^2+4x=0\] \[x^2+8x+16+4x=0\] \[x^2+12x+16=0\]

OpenStudy (anonymous):

you supposed to solve it as well?

OpenStudy (anonymous):

no i got it thank you <3